OpenStudy (anonymous):

Write the equation of a line with a slope of 3 and a y-intercept of 1. y = x + 3 y = 3x + 1 3x + y = 1 <-- I picked this. Am I right? x + 3y = 1

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

It had a slope of -1 but that is the closest out of them all.

pooja195 (pooja195):

y=mx+b m=slope b=y-intercept

OpenStudy (anonymous):

Ah! It must be B than

OpenStudy (anonymous):

Then*

pooja195 (pooja195):

yes :)

OpenStudy (anonymous):

i have one more!

OpenStudy (anonymous):

Should I close this one and tag you in the new one?

pooja195 (pooja195):

just post it here

OpenStudy (anonymous):

Write the equation of the line that passes through (1, 5) and (–2, 14) in slope-intercept form. y = 3x + 2 y = 3x + 8 y = –3x – 2 y = –3x + 8

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] find the slope first

OpenStudy (anonymous):

ok, of one of them right?

pooja195 (pooja195):

Find the slope only first

pooja195 (pooja195):

@Nnesha can u finish plz gotta go

OpenStudy (anonymous):

y = 3x + 2 has a slope of 3 y = 3x + 8 has a slope of 3

OpenStudy (anonymous):

doing the next two one sec

OpenStudy (anonymous):

y = -3x - 2 has a slope of -3

Nnesha (nnesha):

use the formula she gave u plugin the x''s and y's values to find slope (1, 5) and (–2, 14)

OpenStudy (anonymous):

yeah I am. Check the other slopes I already gave out

Nnesha (nnesha):

no we need to find slopes we don't know if it's -3 or positive 3 that's why we should use that formula to find out

Nnesha (nnesha):

\[\LARGE\rm \frac{y_2-y_1}{x_2-x_1}=slope\] \[\huge\rm (x_1,y_1)(x_2 ,y_2)\] \[\huge\rm (1,5)(-2,14)\] x_1 = first x x_2 2nd x